# Systematic Classification of History-Dependent Invariants in the Lorenz System: Regularization Classes and Dynamical Signatures

### v1

@@ Abstract
Building upon the recent discovery of history-dependent dynamical invariants in the Lorenz system, this work presents a systematic exploration of a complete family of such conserved quantities. Through analysis of all 24 permutations in the augmented phase space, we identify eighteen distinct invariant structures ($K_{1}$$K_{18}$) and demonstrate that six permutations yield null results. A master table associates each invariant with its generating permutation, revealing that the class structure emerges naturally from the permutation ordering: Class I ($K_{1}$$K_{6}$, permutations 1abc), Class II ($K_{7}$$K_{12}$, permutations 2abc), and Class III ($K_{13}$$K_{18}$, permutations 3abc). The null permutations 4abc correspond to placing the auxiliary variable first, which yields no valid invariants due to algebraic consistency requirements imposed by Schwarz integrability conditions. A complete regularization procedure is developed for all three classes, with each class characterized by a distinct polynomial factor that removes singularities. Each invariant $K_{n}$ requires its own auxiliary variable $u_{n}(t)$; while the original variable satisfies a linear ODE with $u_{n}$-dependent coefficients, the regularized variable $v_{n}=p\cdot u_{n}$ evolves according to $\dot{v}_{n}=Q_{n}(x,y,z)$, independent of $v_{n}$ itself. Different conservation laws thereby impose distinct measures of trajectory history. Analysis of the functional independence structure establishes that any three invariants selected from different regularization classes provide independent constraints on trajectory histories, with representative invariants from each class defining a canonical “Triad” that characterizes the orbit structure. Statistical analysis of the auxiliary evolution functions $Q_{n}(t)$ reveals that invariants from different regularization classes exhibit distinct dynamical signatures: Class III invariants, whose polynomial part $P_{\mathrm{III}}$ contains terms involving the bilinear product $xy$, display significantly higher intermittency (kurtosis $\kappa_{\mathrm{III}}\approx 15$ versus $\kappa_{\mathrm{I}}\approx 8$ for Class I) and stronger asymmetry (skewness $|S_{\mathrm{III}}|\approx 2.7$ versus $|S_{\mathrm{I}}|\approx 1.6$), correlating with lobe-switching events where the product $xy$ undergoes rapid sign reversal. This geometric sensitivity establishes Class III invariants as probes of topological transitions. The class-dependent divergence of auxiliary evolution functions provides a quantitative measure of how different conservation principles impose incompatible demands on trajectory history, with the variation required to maintain each conservation law encoded in the statistics of the corresponding $\dot{v}_{n}$. High-precision numerical validation confirms conservation of all identified quantities.
@@ I Introduction
The recent demonstration that the Lorenz system possesses history-dependent dynamical invariants has expanded the understanding of conservation laws in dissipative chaotic systems [1]. That work established the existence of a non-local conserved quantity whose constancy along trajectories arises through accumulation of the system’s past via integral terms involving an auxiliary variable. The construction revealed structural organization within an augmented phase space and provided a characterization of unstable periodic orbits as level sets of this conserved functional.
The construction of auxiliary variables to achieve conservation in non-conservative systems connects to several theoretical frameworks. In statistical mechanics, the Mori-Zwanzig formalism demonstrates that projecting high-dimensional dynamics onto lower-dimensional subspaces generates memory kernels, which are convolution integrals over the system’s history [2, 3]. The auxiliary variable $u(t)$ introduced for the Lorenz system represents a deterministic realization of the memory term that emerges when closing an intrinsically dissipative system. This construction aligns with the principles of dimension reduction in dynamical systems.
Independently, Hojman [5] demonstrated that conservation laws can be constructed without recourse to Lagrangian or Hamiltonian formulations, establishing that the connection between symmetries and conserved quantities extends beyond the variational framework. The present construction shares this non-Noetherian character: the invariants $K_{n}$ do not arise from continuous symmetries of an action functional but from algebraic compatibility between the Lorenz flow and the orthogonality ansatz.
@@
However, the constructive method admits multiple distinct choices in the orthogonality ansatz used to augment the three-dimensional Lorenz phase space. The original work noted that “the constructive method employed can yield other conserved quantities,” deferring systematic exploration. This raises specific questions: How many independent history-dependent invariants exist? What determines their mathematical form? Do they possess distinct properties corresponding to different dynamical interpretations?
@@
The present work addresses these questions through systematic analysis. By considering all 24 permutations of the phase space coordinates in the orthogonality construction, we identify exactly eighteen valid invariant structures and demonstrate that six permutations yield null results. The first coordinate in the permutation determines the regularization class. This provides both complete enumeration and structural understanding of the invariant family.
An important insight concerns the nature of the auxiliary variables. Each invariant $K_{n}$ requires its own auxiliary variable $u_{n}(t)$ satisfying a specific evolution equation $\dot{u}_{n}=f_{n}(x,y,z)$. A distinction must be drawn between two formulations: the original auxiliary variable $u_{n}(t)$, which diverges at nullcline crossings where the regularization polynomial vanishes, and the regularized variable $v_{n}(t)=p_{n}(x,y,z)\cdot u_{n}(t)$, which remains bounded for all time. The regularized formulation, in which each invariant takes the form $K_{n}=P_{n}(x,y,z)+v_{n}$, provides the mathematically proper framework. The fluctuation patterns of the regularized variables differ across classes: different invariants exhibit distinct dynamical signatures as they respond to the trajectory’s passage through various regions of phase space. Different conservation laws impose distinct measures of trajectory history. The kinematic variation required to maintain conservation of $K_{n}$ is encoded in the evolution function $Q_{n}(t)=\dot{v}_{n}$, and this variation differs across regularization classes.
The existence of the null class (Theorem 1) demonstrates that not all polynomial structures are compatible with the Lorenz flow: six of 24 permutations fail to yield valid invariants because Schwarz integrability conditions reduce to algebraic contradictions. This construction yields two principal results: the classification of compatible polynomial structures, which encodes information specific to the Lorenz dynamics, and the distinct dynamical signatures exhibited by different regularization classes, which provide complementary probes of attractor geometry.
The invariant $K$ presented in Ref. [1] was derived using an orthogonality ansatz of the form $\nabla C=(\epsilon_{1}\dot{\pi}_{1},\epsilon_{2}\dot{\pi}_{2},\epsilon_{3}\dot{\pi}_{3},\epsilon_{4}\dot{\pi}_{4})$, where each sign $\epsilon_{i}\in\{+1,-1\}$ is assigned independently to each permuted component. This general construction admits $24\times 2^{4}/2=192$ distinct ansätze, of which only 24 correspond to a uniform global sign. The present work focuses on the latter class, which yields the organizational structure described above. Both families share the same regularization polynomials, a connection suggesting underlying structural unity.
We also demonstrate that invariants from different regularization classes exhibit distinct dynamical signatures. Statistical analysis reveals different intermittency patterns in the chaotic regime: Class III invariants, whose polynomial parts $P_{\mathrm{III}}$ contain terms involving the bilinear product $xy$, exhibit enhanced sensitivity to lobe-switching events where this product changes sign. The rapid variation of $xy$ at the separatrix ($x=0$) forces correspondingly rapid variations in $Q_{\mathrm{III}}=-\dot{P}_{\mathrm{III}}$. These events represent topological transitions in the attractor’s symbolic dynamics [6]. The Class III evolution function $Q_{\mathrm{III}}(t)$ functions as a geometric probe that signals each addition to the trajectory’s symbolic sequence, establishing a quantitative link between continuous dynamics and discrete topological structure.
@@
This paper is organized as follows. Section II reviews the constructive orthogonality method and introduces the permutation-based systematic exploration. Section III presents the complete family of eighteen invariants organized by permutation class, including regularization and the proof that permutations beginning with the auxiliary variable yield null results. Section IV establishes the functional independence structure and clarifies the role of invariant-specific auxiliary variables. Section V develops the geometric interpretation of history-dependent constraints. Section VI analyzes transformation properties under the discrete $\mathbb{Z}_{2}$ symmetry. Section VII provides numerical verification and statistical characterization. Section VIII synthesizes the results, including analysis of the physical interpretation of the invariants and their operational meaning.
@@ II.2 Connection to the Mori-Zwanzig Formalism
can be understood as a deterministic resolution of the memory kernel: it explicitly tracks the accumulated effect of the trajectory’s history that would otherwise manifest as non-Markovian dynamics. (The original auxiliary variable $u_{n}=v_{n}/p$ satisfies a more complex linear ODE with state-dependent coefficients; see Sec. III.2.)
A distinction must be drawn between two different approaches to treating unresolved degrees of freedom. In the Optimal Prediction framework of Chorin and collaborators [2], the fluctuating force $F(t)$ is treated stochastically: its effects are averaged over an ensemble, yielding a coarse-grained, probabilistic description. The fluctuation-dissipation theorem [3] then relates the memory kernel to the autocorrelation of this noise, establishing a thermodynamic connection between dissipation and fluctuations.
In contrast, the present construction constitutes a deterministic embedding: by extending the phase space with $v_{n}(t)$, we effectively absorb the orthogonal fluctuations into a resolved variable, achieving an exact closure in which $F(t)\equiv 0$ along any trajectory. This distinction has important consequences. The regularized variable $v_{n}(t)$ in this construction lacks the thermodynamic content of a true memory kernel in the Mori-Zwanzig sense: it does not satisfy a fluctuation-dissipation relation, and its “memory” is purely kinematic rather than thermodynamic. The variable functions as a deterministic history accumulator, recording the trajectory’s past without the statistical properties that would connect it to heat dissipation or entropy production.
This perspective clarifies what the construction achieves: it provides a geometric embedding that renders dissipative dynamics conservative in an extended space, at the cost of introducing a non-local (history-dependent) degree of freedom. The existence of exactly eighteen conserved quantities represents a structural property of the Lorenz equations that is not guaranteed a priori.
@@
Physical scaling. In the original Rayleigh-Bénard convection context from which the Lorenz equations derive [7], the state variables represent amplitudes of truncated Fourier modes: $x$ corresponds to the convective velocity mode, while $y$ and $z$ correspond to temperature perturbation modes. The parameters $\sigma$, $\rho$, and $\beta$ are dimensionless after appropriate scaling by the thermal diffusivity, cell height, and critical Rayleigh number. In this dimensionless formulation, the regularized auxiliary variable $v_{n}(t)$ inherits a kinematic interpretation as a generalized potential that accumulates historical effects of the modal amplitudes. For each regularization class, the polynomial factor $p_{n}(x,y,z)$ combines the physical modes in a specific manner: $p_{\mathrm{I}}=y-x$ couples velocity and temperature perturbations, $p_{\mathrm{II}}=y+x(z-\rho)$ involves the departure from criticality, and $p_{\mathrm{III}}=xy-\beta z$ couples to the nonlinear heat flux term. The invariant $K_{n}=P_{n}+v_{n}$ thereby encodes a balance between instantaneous modal amplitudes and accumulated convective history.
@@ II.3 Permutation-Based Polynomial Construction
and write permutations as four-digit strings. For example, the permutation 1234 corresponds to the identity ordering $(x,y,z,u)$, while 2134 corresponds to $(y,x,z,u)$. When referring to permutation classes, we use the notation nabc where $n\in\{1,2,3,4\}$ is the fixed first index and $a,b,c$ denote the remaining indices in any order.
The inverse construction method. Rather than assuming that the gradient of a conserved quantity equals a permutation of the flow (which would violate Schwarz integrability conditions for most permutations), we employ the permutation as a generator to synthesize a polynomial candidate. Given a permutation $\pi$, we construct a candidate scalar $C$ by summing partial primitives:
@@
where $\pi(\mathbf{f})_{i}$ denotes the $i$-th component of the permuted flow vector and $\mathbf{q}=(x,y,z,u)$. This summation consolidates the contributions from each variable into a coherent algebraic structure, typically yielding a polynomial in $(x,y,z)$ plus terms linear in $u$.
@@
Derivation of auxiliary dynamics. With $C$ constructed, we impose the conservation condition $\dot{C}=0$ and deduce the required auxiliary evolution. The conservation condition yields:
$\frac{\partial C}{\partial x}\dot{x}+\frac{\partial C}{\partial y}\dot{y}+\frac{\partial C}{\partial z}\dot{z}+\frac{\partial C}{\partial u}\dot{u}=0.$ (10)
Solving for $\dot{u}$:
$\dot{u}=-\frac{\frac{\partial C}{\partial x}\dot{x}+\frac{\partial C}{\partial y}\dot{y}+\frac{\partial C}{\partial z}\dot{z}}{\frac{\partial C}{\partial u}}=-\frac{\dot{P}+\dot{p}\cdot u}{p},$ (11)
where we used $\partial C/\partial u=p$. This equation reveals that $\dot{u}$ depends on $u$ itself through the term $(\dot{p}/p)u$. However, when we pass to the regularized variable $v=p\cdot u$, this dependence cancels exactly (Sec. III.2).
The systematic procedure thus consists of:
  1. 1.
    Enumerate all 24 permutations $\pi$ of the indices $\{1,2,3,4\}$.
  2. 2.
    For each permutation, construct the candidate $C$ via Eq. (8).
  3. 3.
    Derive the auxiliary evolution $\dot{u}$ from the conservation condition.
  4. 4.
    Check whether a valid (non-trivial, globally defined) invariant results.
This exhaustive analysis reveals that exactly 18 permutations yield valid invariants ($K_{1}$$K_{18}$), while 6 permutations yield null results. The null permutations are precisely those beginning with the index 4 (the auxiliary variable $u$), for which the integrability conditions fail at the first step (Theorem 1).
Classification by singularity structure. The polynomial $p(x,y,z)$ appearing in the denominator of Eq. (11) determines the singularity structure of the auxiliary evolution. Remarkably, the 18 valid permutations partition into exactly three families based on this polynomial:
@@
This classification by regularization polynomial, rather than by permutation index alone, is the fundamental organizing principle. The correspondence between leading permutation index and class holds for the uniform-sign family explored here, but more general ansätze are classified solely by their regularization polynomial (see Sec. VIII.4).
It should be emphasized that the permutation-based construction provides sufficient conditions for the existence of conserved quantities, not necessary ones. There may exist history-dependent invariants of the Lorenz system that do not arise from any permutation ansatz. The present systematic exploration exhausts the permutation-based family but does not preclude the existence of structurally different invariants constructed by other methods.
II.4 Sign Convention
The invariants $K_{n}$ obtained through direct integration possess an intrinsic global gauge symmetry: if $K$ is a conserved quantity, then K-K is equally conserved. This $\mathbb{Z}_{2}$ symmetry $K\to-K$ reflects the mathematical arbitrariness in the overall sign of any constant of motion, a freedom absent in Noether-derived quantities, which inherit their signs from the underlying Lagrangian structure.
Throughout this work, we present the invariants in their canonical algorithmic form, the expressions that emerge directly from the integration procedure without additional sign manipulation. This choice ensures reproducibility and transparency of the constructive method. For physical applications, one may adopt an alternative positive-definite convention by applying global sign inversion $K_{n}\to-K_{n}$ to individual invariants such that dominant quadratic terms ($y^{2}$, $z^{2}$, $x^{2}$, etc.) appear with positive coefficients.
@@ III.1 Organization by Permutation Class
The null class consists of permutations where the auxiliary variable $u$ occupies the first position. As we shall prove, these permutations cannot yield valid invariants because the resulting consistency equations admit only trivial solutions.
@@ III.2 Regularization Polynomials and Invariant-Specific Auxiliary Variables
Each class is characterized by a distinct regularization polynomial $p(x,y,z)$ that eliminates singularities. Each invariant $K_{n}$ requires its own auxiliary variable $u_{n}(t)$ and regularized variable $v_{n}(t)$. Notation: Throughout this work, lowercase $p$ denotes the regularization polynomial (which multiplies the auxiliary variable), while uppercase $P$ denotes the polynomial part of the invariant (the non-auxiliary component of $K_{n}=P_{n}+v_{n}$).
@@
Note that $v_{n}^{\mathrm{III}}=-p_{\mathrm{III}}u_{n}$; this sign convention is adopted so that the polynomial parts $P_{n}$ of the Class III invariants have leading coefficients consistent with the other classes.
The subscript $n$ on $u_{n}$ and $v_{n}$ is essential: each invariant defines its own auxiliary variable. While all invariants within a class share the same regularization polynomial $p$, they differ in their polynomial parts $P_{n}(x,y,z)$, which determines the evolution equation for each auxiliary variable:
$\dot{v}_{n}=Q_{n}(x,y,z)\equiv-\frac{dP_{n}}{dt}.$ (15)
Since $P_{n}\neq P_{m}$ for $n\neq m$, we have $Q_{n}\neq Q_{m}$, and consequently $u_{n}(t)\neq u_{m}(t)$ along any trajectory.
@@
An important structural feature distinguishes these invariants from standard stochastic treatments: the auxiliary variable $u_{n}(t)$ enters multiplicatively through the coupling $p(x,y,z)\cdot u_{n}$, rather than as an additive noise term. In the Mori-Zwanzig picture, this corresponds to a state-dependent modulation of the memory contribution: the memory kernel’s influence on the dynamics is “gated” by the instantaneous values of the physical coordinates. This multiplicative structure has important consequences: when the polynomial $p(x,y,z)$ vanishes (on nullclines), the memory’s contribution to the invariant is momentarily suppressed, regardless of the accumulated history encoded in $u_{n}$. This gating mechanism may explain the topological robustness observed in the statistical analysis (Sec. VII), as the memory coupling is automatically reduced precisely where the flow undergoes critical transitions.

### v2

# Topological Signatures of History-Dependent Invariants in the Lorenz System: Deterministic Probes of Lobe Dynamics

@@ Abstract
Building upon the recent discovery of history-dependent dynamical invariants in the Lorenz system, this work presents a systematic classification of a complete family of such conserved quantities and demonstrates their utility as deterministic probes of topological transitions. Through analysis of all 24 permutations in the augmented phase space, we identify eighteen distinct invariants ($K_{1}$$K_{18}$) organized into three regularization classes, while six permutations yield null results due to Schwarz integrability constraints. Each class is characterized by a distinct regularization polynomial that vanishes on the corresponding nullcline of the Lorenz flow: $p_{\mathrm{I}}=y-x$, $p_{\mathrm{II}}=y+x(z-\rho)$, and $p_{\mathrm{III}}=xy-\beta z$.
The central result concerns the dynamical signatures that distinguish regularization classes. Statistical analysis of $N=4\times 10^{6}$ trajectory samples reveals that Class III invariants, whose polynomial parts contain the bilinear product $xy$, function as geometric probes of lobe-switching events. In the chaotic regime ($\rho=28$), Class III exhibits elevated kurtosis ($\kappa_{\mathrm{III}}\approx 15.8$ versus $\kappa_{\mathrm{I}}\approx 8.2$) with heavy tails extending beyond $\pm 5\sigma$, while both classes show pronounced negative skewness ($S_{\mathrm{I}}\approx-1.66$, $S_{\mathrm{III}}\approx-2.73$). This divergence is absent in the pre-chaotic regime ($\rho=23$), where both classes yield statistically similar distributions ($\kappa\approx 28.2$), confirming that the signatures arise specifically from the interaction between polynomial structure and chaotic dynamics.
Remarkably, the differential response $\Delta S=Q_{\mathrm{III}}-Q_{\mathrm{I}}$ between Class III and Class I evolution functions provides advance warning of separatrix crossings, with a prediction horizon governed by a shifted power law $\Delta t=t_{\min}+C\mathcal{A}^{-n}$. Parametric analysis across $\beta\in[2.0,3.2]$ reveals a robust linear relationship $n=m\beta+b$ with slope $m=0.196\pm 0.02$ ($R^{2}=0.94$), establishing that the predictive exponent scales with the stable eigenvalue. At the standard value $\beta=8/3$, the fitted exponent $n\approx 1.0$ reflects the characteristic relaxation timescale along the stable manifold. This connection between algebraic classification and predictive capability demonstrates that Class III invariants encode the symbolic dynamics of the attractor’s template structure. High-precision numerical validation using 80-digit arithmetic confirms conservation of all identified quantities to $O(10^{-36})$.
@@ I Introduction
The recent demonstration that the Lorenz system possesses history-dependent dynamical invariants has expanded the understanding of conservation laws in dissipative chaotic systems [1]. That work established the existence of a non-local conserved quantity whose constancy along trajectories emerges through accumulation of the system’s past via integral terms involving an auxiliary variable. The construction revealed structural organization within an augmented phase space and provided a characterization of unstable periodic orbits as level sets of this conserved functional.
The physical significance of these invariants extends beyond mathematical curiosity. In the Mori-Zwanzig formalism [2, 3], memory effects arise naturally when projecting high-dimensional dynamics onto a reduced set of observables. The auxiliary variable $u(t)$ in this construction plays an analogous role: it accumulates the system’s history in a manner that compensates for the apparent dissipation in the original three-dimensional flow. The regularization polynomials $p(x,y,z)$ that characterize each class are not arbitrary mathematical artifacts; they are precisely the nullcline conditions of the Lorenz equations, suggesting that the invariant structure is intimately connected to the geometric skeleton of the flow.
@@
The constructive method admits multiple distinct choices in the orthogonality ansatz used to augment the three-dimensional Lorenz phase space. The original work noted that “the constructive method employed can yield other conserved quantities,” deferring systematic exploration. This raises specific questions: How many independent history-dependent invariants exist? What governs their mathematical form? Do they possess distinct properties corresponding to different dynamical interpretations?
@@
The present work addresses these questions through systematic analysis and reveals an unexpected connection between algebraic structure and dynamical sensitivity. By considering all 24 permutations of the phase space coordinates in the orthogonality construction, we identify exactly eighteen valid invariant structures and demonstrate that six permutations yield null results. The first coordinate in the permutation determines the regularization class, which in turn determines how the invariant responds to chaotic dynamics.
The central finding is that Class III invariants function as deterministic probes of topological transitions. The polynomial parts $P_{\mathrm{III}}$ contain terms involving the bilinear product $xy$, which changes sign at every lobe-switching event where the trajectory crosses the separatrix $x=0$. Statistical analysis reveals that Class III evolution functions exhibit large-amplitude spikes synchronized with these crossings, while Class I evolution functions vary more smoothly. This differential response enables construction of a topological precursor signal that provides advance warning of separatrix crossings, with predictive capability governed by a power law whose exponent scales linearly with the stable eigenvalue $\beta$.
Independently, Hojman [5] demonstrated that conservation laws can be constructed without recourse to Lagrangian or Hamiltonian formulations, establishing that the connection between symmetries and conserved quantities extends beyond the variational framework. The present construction shares this non-Noetherian character: the invariants $K_{n}$ do not arise from continuous symmetries of an action functional but emerge from algebraic compatibility between the Lorenz flow and the orthogonality ansatz.
@@
This paper is organized as follows. Section II reviews the constructive orthogonality method and introduces the permutation-based systematic exploration. Section III presents the complete family of eighteen invariants organized by permutation class, including regularization and the proof that permutations beginning with the auxiliary variable yield null results. Section IV establishes the functional independence structure and clarifies the role of invariant-specific auxiliary variables. Section V develops the geometric interpretation of history-dependent constraints. Section VI analyzes transformation properties under the discrete $\mathbb{Z}_{2}$ symmetry. Section VII provides numerical verification, statistical characterization, and the analysis of topological precursor detection. Section VIII synthesizes the results, including analysis of the physical interpretation of the invariants and their operational meaning.
@@ II.2 Connection to the Mori-Zwanzig Formalism
can be understood as a deterministic resolution of the memory kernel: it explicitly tracks the accumulated effect of the trajectory’s history that would otherwise manifest as non-Markovian dynamics.
A distinction must be drawn between two different approaches to treating unresolved degrees of freedom. In the Optimal Prediction framework of Chorin and collaborators [2], the fluctuating force $F(t)$ is treated stochastically: its effects are averaged over an ensemble, yielding a coarse-grained, probabilistic description. In contrast, the present construction constitutes a deterministic embedding: by extending the phase space with $v_{n}(t)$, we effectively absorb the orthogonal fluctuations into a resolved variable, achieving an exact closure in which $F(t)\equiv 0$ along any trajectory.
@@
Physical scaling. In the original Rayleigh-Bénard convection context from which the Lorenz equations derive [7], the state variables represent amplitudes of truncated Fourier modes: $x$ corresponds to the convective velocity mode, while $y$ and $z$ correspond to temperature perturbation modes. The parameters $\sigma$, $\rho$, and $\beta$ are dimensionless after appropriate scaling by the thermal diffusivity, cell height, and critical Rayleigh number. In this dimensionless formulation, the regularized auxiliary variable $v_{n}(t)$ inherits a kinematic interpretation as a generalized potential that accumulates historical effects of the modal amplitudes. For each regularization class, the polynomial factor $p_{n}(x,y,z)$ combines the physical modes in a specific manner: $p_{\mathrm{I}}=y-x$ couples velocity and temperature perturbations, $p_{\mathrm{II}}=y+x(z-\rho)$ involves the departure from criticality, and $p_{\mathrm{III}}=xy-\beta z$ couples to the nonlinear heat flux term.
@@ II.3 Permutation-Based Polynomial Construction
and write permutations as four-digit strings. For example, the permutation 1234 corresponds to the identity ordering $(x,y,z,u)$, while 2134 corresponds to $(y,x,z,u)$.
The inverse construction method. Rather than assuming that the gradient of a conserved quantity equals a permutation of the flow, we employ the permutation as a generator to synthesize a polynomial candidate. Given a permutation $\pi$, we construct a candidate scalar $C$ by summing partial primitives:
@@
where $\pi(\mathbf{f})_{i}$ denotes the $i$-th component of the permuted flow vector and $\mathbf{q}=(x,y,z,u)$.
@@
Classification by singularity structure. The polynomial $p(x,y,z)$ appearing in the denominator of the auxiliary evolution determines the singularity structure. Remarkably, the 18 valid permutations partition into exactly three families based on this polynomial:
@@
This classification by regularization polynomial is the fundamental organizing principle.
@@ III.1 Organization by Permutation Class
The null class consists of permutations where the auxiliary variable $u$ occupies the first position. These permutations cannot yield valid invariants because the resulting consistency equations admit only trivial solutions.
@@ III.2 Regularization Polynomials and Invariant-Specific Auxiliary Variables
Each class is characterized by a distinct regularization polynomial $p(x,y,z)$ that eliminates singularities. Each invariant $K_{n}$ requires its own auxiliary variable $u_{n}(t)$ and regularized variable $v_{n}(t)$.
@@
It is interesting to note that the multiplicative structure $p(x,y,z)\cdot u_{n}$ implies a “gating” mechanism: when the polynomial $p(x,y,z)$ vanishes on nullclines, the memory’s contribution to the invariant is momentarily suppressed, regardless of the accumulated history encoded in $u_{n}$. This gating mechanism may explain the topological robustness observed in the statistical analysis, as the memory coupling is automatically reduced precisely where the flow undergoes critical transitions.